OpenStudy (anonymous):

solve the inequality, then write solution set in interval notation 7x+2+x(greater than or equal to) 1 +3x-19

OpenStudy (anonymous):

who is the guy next to you in the pic

OpenStudy (anonymous):

\[7x+2+x \ge1+3x-19\rightarrow8x+2\ge3x-18\rightarrow5x \ge-20\rightarrow x \ge-5 or [-5,\infty)\]

OpenStudy (anonymous):

\[[-5,\infty)\]

OpenStudy (anonymous):

the answer has only -4 choices...

OpenStudy (anonymous):

which one are in you in the pic

OpenStudy (anonymous):

i mean who is the guy

OpenStudy (anonymous):

my boyfriend

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

do you have a bigger pic

OpenStudy (anonymous):

you dont even have a pic..im just here to get help with math...

OpenStudy (anonymous):

i can see you there

OpenStudy (anonymous):

with my eagle eyes

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

whats up ?

OpenStudy (anonymous):

i can see you in the pic :O

OpenStudy (anonymous):

solve the inequality. 5<1-5m<8

OpenStudy (anonymous):

\[4<-5m<7\rightarrow-4/5<m<-7/5\]

OpenStudy (anonymous):

\[6<2x+4\le12\]

OpenStudy (anonymous):

\[2<2x \le8\rightarrow1<x \le4\]

OpenStudy (anonymous):

but what is the interval notation and how do i graph it?

OpenStudy (anonymous):

how do i write as a compound inequality?

OpenStudy (anonymous):

interval notation are: (-4/5,-7/5) and for the last inequality is (1,4].

OpenStudy (anonymous):

thats not even one of the choices for 1 and 4 though..? theres (1,4) [1,4] (-4,-1) [-4,-1}